OpenStudy (anonymous):

the machine gun fires 25 gram bullets at the rate of 600 bullets per minute with a speed of 200m/s. calculate the force requred to keep the gun in position.

OpenStudy (anonymous):

plzz do....

OpenStudy (anonymous):

.025*600*60*200

OpenStudy (anonymous):

force is the rate of change of momentum

OpenStudy (anonymous):

does the number of bullets being shot being taken into account? why not just use 25g?

OpenStudy (anonymous):

is the ans is 5000N??

OpenStudy (anonymous):

yea number of bullets shud be considered

OpenStudy (anonymous):

means the ans is 5000N ??

OpenStudy (anonymous):

????????????

OpenStudy (anonymous):

calculate it.......is the ans 5000?

OpenStudy (anonymous):

but F = (mVf - mVi) / t ...the initial velocity is 0??

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

is the ans is right?

OpenStudy (anonymous):

i think it's 180000 !!!

OpenStudy (anonymous):

haha! I got that too.

OpenStudy (anonymous):

isnt it was 5000 as, (mass/time)* velocity??

OpenStudy (anonymous):

F = ma

OpenStudy (anonymous):

well do it this way calculate impluse it is mdv I=F dt u l get dt.....

OpenStudy (anonymous):

didn't get u !!

OpenStudy (anonymous):

but in this case, it's related to momentum.

OpenStudy (anonymous):

yea....so? u need force r8 um @kulveen do u kno wat impulse is?

OpenStudy (anonymous):

yaa bt f is the rate of change of momentum!!

OpenStudy (anonymous):

ok...calculate the rate of change.............. mdv/dt...........

OpenStudy (anonymous):

oh, I was saying it's easier to use momentum to find rather than F=ma, after your explanation..

OpenStudy (anonymous):

:/

OpenStudy (anonymous):

no i dont meaning of impulse... :P

OpenStudy (anonymous):

how about momentum?

OpenStudy (anonymous):

fine then do it this way..........for a single bullet the change of momentum is mdv....... nd dt wud be 1/600/60..................

OpenStudy (anonymous):

wait....

OpenStudy (anonymous):

cant we do in this way- Force is the rate of change of momentum = (mass/time)*velocity

OpenStudy (anonymous):

ok....u can

OpenStudy (anonymous):

can u do it in this way and tl me the ans?

OpenStudy (anonymous):

wat's m/t = -25/(1/600/60) nd v = 200

OpenStudy (anonymous):

uh, I got 50N.

OpenStudy (anonymous):

that's 0.025

OpenStudy (anonymous):

no u wud be in KN...........it wont be that less

OpenStudy (anonymous):

is this the rght ans?? 0.025??

OpenStudy (anonymous):

noo lol ans is 180KN

OpenStudy (anonymous):

tl me the right ans....

OpenStudy (anonymous):

u got options?

OpenStudy (anonymous):

I agree....because you don't have the exact time...if you use F =mass/time *velocity it will be inaccurate in my opinion, sorry! I forgot about the conversion :P anwer : 180kN

OpenStudy (anonymous):

noooo

OpenStudy (anonymous):

ans is 180KN

OpenStudy (anonymous):

can u do it? here

OpenStudy (anonymous):

it did it r8..........

OpenStudy (anonymous):

means thoroly can u do it heree??

OpenStudy (anonymous):

.025*600*60*200 :/